OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2) (x +7)=0 A. 2, -7 B. 2, 7 C. –1,1 D. –2, -7

OpenStudy (anonymous):

it means that either (x - 2) = 0 or (x + 7) = 0. only one at a time being necessary since anything times zero is zero. so we have x - 2 = 0 and x + 7 = 0 which answer do u think?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

a is right

OpenStudy (johnweldon1993):

Well lets check 'b' (2 , 7) your 2 equations are x - 2 = 0 x = 2 ...good here x + 7 = 0 14 = 0....not good here... Yes ....A is correct

OpenStudy (anonymous):

ty